# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ II.1.4 The linear-zigzag transition
These conjectures are put on a solid ground by a statistical mechanics approach [109], that derives the Landau theory of the linear-zigzag instability starting from the classical potential energy of an ion chain. This theory shows that the linear-zigzag transition in ion Coulomb chains is of the same universality class of the mean-field phase transition from paramagnet to ferromagnet. Its predictions were so far confirmed by experimental studies, as we detail in Sec. III.4. In what follows we review the main arguments and steps of the derivation in Ref. [109].
Figure 13: Left panel: Summary of the experimentally observed configuration as a function of the potential depth and of the linear ion density. These two parameters can be combined to give the normalized linear particle density $\lambda$, which fully determines the ion configuration (see text). The observed configurations are labeled by different symbols for each structure. The straight lines show critical $\lambda$ values separating the regions of different theoretically expected structures, see [136]. Right panel: Color-coded images of crystalline structures of laser-cooled 24Mg+ ions in a quadrupole trap. Adapted from [22].
Figure 14: Normal modes of a uniform chain as a function of the quasimomentum $k$ for (left) the linear array, $\nu_{t}>\nu_{t}^{(c)}$, (middle) at the transition, $\nu_{t}=\nu_{t}^{(c)}$, and (right) a zigzag array, $\nu_{t}<\nu_{t}^{(c)}$. Left and central panels: The red dashed line denotes the the transverse vibrations, Eq. (41). Central panel: At the critical frequency, the mode at vanishing frequency is the zigzag mode. Right panel: The Brillouin zone of the zigzag array is half the size of the linear chain, due to the period doubling of the zigzag structure. Adapted from Ref. 109.
@@
where $\Theta_{k}^{(s)}$ are the Fourier transform of the axial displacements $q_{j}$ and $\Psi^{y(s)}_{k},\Psi^{z(s)}$ of the transverse displacements $y_{j}$ and $z_{j}$, respectively, and $s=\pm$ denotes the $\cos$ and $\sin$ standing wave. The left panel of Fig. 14 displays the normal mode spectrum. For the linear-zigzag instability, the dispersion relation of the transverse excitation is crucial. Its explicit form [134, 109],
$\beta(k)^{2}=\nu_{t}^{2}-K_{0}\left(\frac{L}{a}\right)^{3}\sum_{j=1}^{N/2}\frac{1}{j^{3}}\sin^{2}\frac{jka}{2}$ (41)
allows one to extract several key observations: (i) the minimal transverse frequency $\omega^{\perp}(k)$ is at the quasimomentum $k_{0}$. These are the eigenmodes with the shortest wavelength, $2a$, also known as zigzag eigenmodes because neighboring ions oscillate with opposite phase, $y_{i}=(-1)^{i}\Phi_{k_{0}}^{(y)}$ and $z_{i}=(-1)^{i}\Phi_{k_{0}}^{(z)}$. (ii) The linear chain becomes unstable at $\nu_{t}=\nu_{t}^{(c)}$ such that the eigenfrequency of the zigzag mode vanishes, $\beta(k_{0})=0$ (see central panel of Fig. 14). The critical transverse trap frequency takes the explicit form
@@
with $\zeta(j)$ Riemann’s zeta function and $\omega_{0}=\sqrt{{Q^{2}}/(4\pi\epsilon_{0}ma^{3})}$. In terms of this notation, the normalized particle density of Fig. 13 becomes $\lambda=(3\omega_{0}^{2}/\nu_{t}^{2})^{1/3}$, so that its critical value for linear-zigzag transition is $\lambda_{c}=0.709=(3/7\zeta(3))^{1/3}$.
@@
$V_{\rm 0}(\varrho)=\frac{1}{2}m\left(\nu_{t}^{2}-\nu_{t}^{(c)2}\right)\varrho^{2}+{\mathcal{A}}\varrho^{4}$ (43)
@@
Figure 15: Transverse equilibrium displacement $b=\varrho$, in units of the interparticle spacing $a$, as a function of the transverse frequency $\nu_{t}$ in units of $\nu_{t}^{(c)}$. On the right of the curve the ion crystal is a linear chain. In the region on the left of the curve it exhibits a zigzag structure. Adapted from Ref. 109.
This derivation is straighforwardly extended to geometries where the transverse confining potential is anisotropic, such that the confining frequency $\nu_{t}$ appearing in Eq. (27) is replaced by two different values, $\nu_{t,y}$ and $\nu_{t,z}$. In this case the anisotropy pins the zigzag orientation and aligns it in the direction of the weakest transverse confinement. For, e.g., $\nu_{t,y}<\nu_{t,z}$, Eq. (44) gives the displacement $\varrho=\Psi_{k_{0}}^{(y)}$ and the angle, determining the plane, takes the two possible values $\varphi=0,\pi$: the Landau potential $V_{0}$, Eq. (43), now possesses two degenerate minima at $\phi_{j}=\pm\bar{\varrho}$ with $\bar{\varrho}$ given by Eq. (44). Thus, the anisotropy gaps out the Goldstone mode described above and the zigzag structure breaks the symmetry by reflection about the axis. This reduces the symmetry of the transition from XY to Ising symmetry. The potential close to the transition, including the first non-vanishing order in the small parameter of the gradient expansion $|\delta k|^{2}a^{2}$, takes the form [109, 294, 293]:
$V[\{\phi_{j}\}]=\sum_{j=1}^{N}V_{0}(\phi_{j})+\frac{1}{2}K\sum_{j=1}^{N}\left(\phi_{j}-\phi_{j+1}\right)^{2}$ (45)
@@
This model is based on several conditions: critical behavior is strictly found for $N\to\infty$ and temperatures $T=0$. In a linear Paul trap, where the density is inhomogeneous, the zigzag structure forms first at the chain center, where the density is maximal [278]. For small chains the zigzag structure displaces the ions not only in the transverse but also in the axial direction. This effect becomes negligible close to the transition points for chains of dozen of ions [109]. For chains of this size, the zigzag eigenmodes are localized at the chain center and are well approximated by standing waves as in a uniform chain  [^1] Interestingly, at the trap center Eq. (38) reduces to a wave equation in the thermodynamic limit, $I[w(\xi)]\sim\log N\frac{\partial^{2}}{\partial\xi^{2}}w(\xi)$ and the eigenmodes at the center are phononic waves. This expression also reproduces the behavior of the axial dispersion relation of a periodic chain for $k\sim\pi/(Na)$, and $\omega^{\|}\sim k\sqrt{|\log k|}$ [7].[224]. In [167] it was experimentally shown that the micromotion modifies the eigenfrequencies close to the transition [167]. The measurements are in agreement with a theoretical treatment, which includes the periodic drive of the Paul trap [181] (see also Sec. III.1).
@@
Figure 16: (a) Low-frequency spectrum as a function of the aspect ratio $\alpha$. The solid lines are the modes of the harmonic crystal at T = 0, the symbols refer to the experimental measurements obtained with different techniques along with estimated uncertainties. The ions are laser cooled close to the Doppler limit. Comparison with molecular dynamics simulations gives $T\approx 3.5$ mK. (b) Schematic illustration of the Landau potential in the linear ($\alpha>\alpha_{c}$) and in the zigzag symmetry-broken phase ($\alpha<\alpha_{c}$). The coordinate $y$ indicates the transverse displacement of the central ion from the chain axis. Thermal excitations can overcome the barrier between the two degenerate zigzag configuration by inducing collective jumps of the crystal configuration. (c) Average dwelling time $\tau_{e}$, calculated using Kramers formula, in units of the zigzag mode period $T_{zz}=1/\omega_{k_{0}}$. In the simulations, temperatures $T$ are 0.1 mK (blue circles), 0.5 mK (orange squares), 2.0 mK (green diamonds), 3.5 mK (red triangles). The lines are a guide for the eye. For T=T=0.1 mK and 0.5 mK, missing points indicate no switches were observed during the simulation time. The dotted horizontal line indicates $\tau_{e}=T_{zz}$ . Inset shows the potential barrier $E_{B}$ for different trapping ratios $\alpha$. Adapted from Ref. 174.
@@
At finite, non-vanishing temperatures, the theory predicts that the linear-zigzag transition is a smooth crossover. For finite chains, nevertheless, a relatively sharp transition can be measured when the chain is prepared at temperatures below a critical value $T_{c}$ that depends on the number of ions $N$. In Ref. [174] the normal-mode spectrum of a trapped ion chain was measured at the symmetry-breaking linear to zigzag transition and at finite temperatures. Spectroscopy was performed by modulating the amplitude of the Doppler cooling laser to excite and measure mode oscillations of a chain composed of 30 172Yb ions. The expected mode softening at the critical point, a signature of the second-order transition, was not observed, see the left panel of Fig. 16. The spectroscopic signal at the transition is reproduced by molecular dynamics simulations, demonstrating that the disappearance of the mode softening is due to the finite temperature of the chain. The experimental low-frequency spectrum about the transition point is also reproduced by an effective theory. The effective model starts from the Landau model and includes terms of the Taylor expansion which couple the soft mode with high-frequency modes, with a coupling strength that scales with the temperature. Inspection of the numerical trajectories shows that the ions collectively jump between the two ground-state configurations of the symmetry-broken phase, as illustrated in the right panel of Fig. 16. This jump rate is qualitatively reproduced using Kramers’ formalism [76]. In this regime, where the ion dynamics consists of thermally activated jumps between the two zigzag configurations, a time-averaged measurement of the mean transverse displacement detects a linear chain where one would instead expect a zigzag structure. The transition to a zigzag is measured at trap aspect ratios below the critical one, resembling a temperature-induced shift of the transition point [125, 76, 192].
@@
In Ref. [65] a quantum thermodynamic study of the linear-zigzag transition was performed. Tt is argued that the average work distribution diverges at the linear-zigzag critical point. It is proposed to reveal this prediction by performing small quenches of the ion trap frequency across the transition and measuring the ion transverse displacements distribution. Interferometric protocols, based on spin-motion entanglement across the linear-zigzag instability, could access the autocorrelation function of the crystal [69] and measure the normal mode spectrum at criticality [12, 13]. However, the interferometric signals are relatively sensitive to thermal effects, which tend to suppress entanglement; an assessment is found in Ref. [14]. So far, the universal critical exponents of Landau theory have been indirectly verified by experiments analysing the scaling of defects by slowly quenching the trap aspect ratio across the linear-zigzag phase transition, see Sec. III.4. In Ref. [339] sideband spectroscopy was performed at the linear-zigzag transition, revealing quantum effects. This suggests a potential access to a structural transition driven by quantum fluctuations, that is discussed in the next subsection.
@@
Figure 17: Left panel. Top: Sample images of the two symmetry broken equilibrium structures of the five 171Yb ion crystal far from the critical point. Bottom: Simulated quantum energy-level spectrum for the first two excited states of the transverse-y zigzag mode of a four-ion crystal, calculated relative to the ground state. Color shading of the levels indicates the strength of Raman sideband coupling to the ground state for a central ion (see sidebar scale). Also shown is the classical small-oscillation prediction and the form of the zigzag potential on either side of the transition. Right panel: measured frequency of first and second upper Raman sidebands for the transverse-y zigzag mode as a function of secular trap frequency $\nu_{y}$ relative to critical value $\nu_{y,c}=760$ kHz. Insets show the full range of data acquisition and a close-in view near the critical point. Solid lines are quantum energy-level differences of $n=1$ and $n=2$ number states with respect to the $n=0$ ground state for a quartic potential. Line shading corresponds to Raman coupling strength (see side-bar scale). Green dotted line shows the classical small-oscillation prediction for the full pseudopotential. From 339.
@@ II.1.5 The quantum linear-zigzag transition
Structural phase transitions in ion Coulomb crystals are typically modelled by the configurations that minimize the potential energy. This approach, despite the fact it discards the kinetic energy, proves to be sufficiently accurate to reproduce the experimental results for Doppler-cooled chains. It works even down to low dimensions, thanks to the long-range nature of the Coulomb interactions. For this reason, the predictions of Landau theory for the linear-zigzag transition are confirmed by several experimental results within the level of accuracy permitted by the temperature and by the measurement apparatus. The possibility to cool relatively long chains close to the zero-point motion, on the other hand, paves the way for verifying the effects of quantum fluctuations at the linear-zigzag transition. In the near vicinity of the transition, in fact, the competing influences of the Coulomb interaction and the confining potential balance each other out, and fluctuations are expected to become significant. In Ref. [262] tunnelling across the linear-zigzag transition was discussed for an anisotropic model, which confines the motion of the ions on a plane. By means of a field-theoretical model, which was postulated on the basis of plausible assumptions, the tunnelling rate between the two symmetry-broken zigzag ground states were estimated for chains of few ions. In Ref. [339] it was possible to experimentally reveal quantum tunnelling between zigzag configurations in a chain composed of three to five 171Yb ions and confined by a linear Paul trap. Figure 17 displays the spectroscopic signal of a Raman transition resolving the anharmonic spectrum of excitations of the zigzag mode, demonstrating the effect of quantum dynamics.
For large chains, where the linear-zigzag transition is classically a Landau second-order phase transition, quantum effects modify the critical behavior. In Refs. [294, 293] a quantum field theoretical model was derived for an anisotropic potential as in Ref. [339], which pins the zigzag orientation in the xyx-y plane. The procedure permits to map the quantum linear-zigzag transition to an Ising model in transverse field. The derivation incorporates quantum fluctuations in the classical model, Eq. (45), and leads to a 1+1 dimensional quantum field theory [294] described by the partition function $Z=\int\mathcal{D}\phi e^{-S[\phi]/\hbar}$ with the Euclidean action
@@
$+\left.\frac{1}{2}K(\phi_{j}-\phi_{j+1})^{2}\right]$ (46)
@@
$H_{\mathrm{I}}=-\sum_{j=1}^{N}\left(J\hat{\sigma}_{j}^{z}\hat{\sigma}_{j+1}^{z}+h\hat{\sigma}_{j}^{x}\right)$ (47)
@@
$J=K\bar{\varrho}^{2}=K\frac{m}{4\mathcal{A}}\left(\nu_{t}^{(c)2}-\nu_{t}^{2}\right)$ (48)
@@
Figure 18: Phase diagram for a linear-zigzag transition, according to the mapping to the quantum Ising chain, where $T$ is the temperature and $\varepsilon$ is defined in (49). The quantum critical point, at $\varepsilon_{c}>0$, separates the linear from the zigzag phase at $T=0$. For $0<\varepsilon<\varepsilon_{c}$ quantum fluctuations dominate, and the crystal is in the linear (disordered) phase. The dashed lines indicate the boundaries of the quantum critical region. From 294.
Figure 19: Computed critical exponents and central charge of the action in Eq. (46). The theory values are the ones of the Ising model in transverse field, see e.g. 272. From 296.
@@
$\tilde{\hbar}=\frac{\hbar}{\sqrt{maQ^{2}\log 2}}\,.$ (50)
where the dimensionless variable $\tilde{\hbar}$ is the effective Planck constant. For the typical parameters of ion chains $\varepsilon$ takes values ranging between $10^{-5}$ and $10^{-4}$, see [296].
In Ref. [251] the shift $\varepsilon$ was computed as a function of the small dimensionless parameter, $\tilde{\hbar}$. This was perfomed by rewriting the partition function $Z=\int{\mathcal{D}}\phi e^{-S[\phi]/\hbar}$ in the form

### v2

@@ II.1.4 The linear-zigzag transition
We devote the rest of this section to the structural instability of the linear chain. The demonstration that this is a second order phase transition, of the Ising universality class of magnetic systems, has permitted to directly connect the physics of ion Coulomb crystals with statistical mechanics and condensed matter concepts and models. This has paved the way to a series of theoretical and experimental investigations on critical properties which we review in what follows. It has further inspired and set the basis for subsequent works on the formation and stability of topological defects, see Secs. III.3-III.6.
@@
These conjectures are put on a solid ground by a statistical mechanics approach, that derives the Landau theory of the linear-zigzag instability starting from the classical potential energy of an ion chain [118]. This theory shows that the linear-zigzag transition in ion Coulomb chains is of the same universality class of the mean-field phase transition from paramagnet to ferromagnet. Its predictions were so far confirmed by experimental studies, as we detail in Sec. III.4. In what follows we review the main arguments and steps of the derivation in Ref. [118].
Figure 13: Left panel: Summary of the experimentally observed configuration as a function of the potential depth and of the linear ion density. These two parameters can be combined to give the normalized linear particle density $\lambda$, which fully determines the ion configuration (see text). The observed configurations are labeled by different symbols for each structure. The straight lines show critical $\lambda$ values separating the regions of different theoretically expected structures, see [145]. Right panel: images of crystalline structures of laser-cooled 24Mg+ ions in a quadrupole trap. Adapted from [25].
Figure 14: (Color online) Normal modes of a uniform chain as a function of the quasimomentum $k$ for (left) the linear array, $\nu_{t}>\nu_{t}^{(c)}$, (middle) at the transition, $\nu_{t}=\nu_{t}^{(c)}$, and (right) a zigzag array, $\nu_{t}<\nu_{t}^{(c)}$. Left and central panels: The red dashed line denotes the the transverse vibrations, Eq. (41). Central panel: At the critical frequency, the mode at vanishing frequency is the zigzag mode. Right panel: The Brillouin zone of the zigzag array is half the size of the linear chain, due to the period doubling of the zigzag structure. Adapted from Ref. 118.
@@
where $\Theta_{k}^{(s)}$ are the Fourier transform of the axial displacements $q_{j}$ and $\Psi^{y(s)}_{k},\Psi^{z(s)}$ of the transverse displacements $y_{j}$ and $z_{j}$, respectively, while $s=\pm$ denotes the $\cos$- and $\sin$- standing waves. The left panel of Fig. 14 displays the normal mode spectrum. For the linear-zigzag instability, the dispersion relation of the transverse excitation is crucial. Its explicit form [146, 118],
$\beta(k)^{2}=\nu_{t}^{2}-K_{0}\left(\frac{L}{a}\right)^{3}\sum_{j=1}^{N/2}\frac{1}{j^{3}}\sin^{2}\frac{jka}{2}\,$, (41)
allows one to extract several key observations: (i) the minimal transverse frequency $\omega^{\perp}(k)$ is at the quasimomentum $k_{0}=\pi/a$. These are the eigenmodes with the shortest wavelength, $2a$, also known as zigzag eigenmodes because neighboring ions oscillate with opposite phase, $y_{i}=(-1)^{i}\Phi_{k_{0}}^{(y)}$ and $z_{i}=(-1)^{i}\Phi_{k_{0}}^{(z)}$. (ii) The linear chain becomes unstable when the eigenfrequency of the zigzag mode vanishes, $\beta(k_{0})=0$ (see central panel of Fig. 14). This determines the critical transverse trap frequency
@@
with $\zeta(j)$ Riemann’s zeta function and $\omega_{0}=\sqrt{{Q^{2}}/(4\pi\epsilon_{0}ma^{3})}$. In terms of this notation, the normalized particle density of Fig. 13 becomes $\lambda=(3\omega_{0}^{2}/\nu_{t}^{2})^{1/3}$, so that its critical value for linear-zigzag transition is $\lambda_{c}=(3/7\zeta(3))^{1/3}\approx 0.709$.
@@
$V_{\rm 0}(\varrho)=\frac{1}{2}m\left(\nu_{t}^{2}-\nu_{t}^{(c)2}\right)\varrho^{2}+{\mathcal{A}}\varrho^{4}\,$, (43)
@@
Figure 15: Transverse equilibrium displacement $b=\varrho$, in units of the interparticle spacing $a$, as a function of the transverse frequency $\nu_{t}$, in units of $\nu_{t}^{(c)}$. On the right of the curve the ion crystal is a linear chain. In the region on the left of the curve it exhibits a zigzag structure. Adapted from Ref. 118.
This derivation is straighforwardly extended to geometries where the transverse confining potential is anisotropic, such that the confining frequency $\nu_{t}$ appearing in Eq. (27) is replaced by two different values, $\nu_{t,y}$ and $\nu_{t,z}$. In this case the anisotropy pins the zigzag orientation and aligns it in the direction of the weakest transverse confinement. For, e.g., $\nu_{t,y}<\nu_{t,z}$, Eq. (44) gives the displacement $\varrho=\Psi_{k_{0}}^{(y)}$ and the angle, determining the plane, takes the two possible values $\varphi=0,\pi$: the Landau potential $V_{0}$, Eq. (43), now possesses two degenerate minima at $\phi_{j}=\pm\bar{\varrho}$ with $\bar{\varrho}$ given by Eq. (44). Thus, the anisotropy favors a zigzag structure confined to a particular plane. This zigzag configuration spontaneously breaks the reflection symmetry about the axis. As a consequence, the continuous XY symmetry of the transition is reduced to a discrete Ising symmetry. The potential close to the transition, including the first non-vanishing order in the small parameter of the gradient expansion $|\delta k|^{2}a^{2}$, takes the form [328]:
$V[\{\phi_{j}\}]=\sum_{j=1}^{N}V_{0}(\phi_{j})+\frac{1}{2}K\sum_{j=1}^{N}\left(\phi_{j}-\phi_{j+1}\right)^{2}\,$, (45)
@@
The Landau model is based on several conditions: critical behavior is strictly found for $N\to\infty$ and temperatures $T=0$. In a linear Paul trap, where the density is inhomogeneous, the zigzag structure forms first at the chain center, where the density is maximal [311]. For small chains the zigzag structure displaces the ions not only in the transverse but also in the axial direction. This effect becomes negligible close to the transition in chains of dozen of ions [118], where the zigzag eigenmodes are localized at the chain center and are well approximated by phononic waves  [^1] At the trap center Eq. (38) reduces to a wave equation in the thermodynamic limit, $I[w(\xi)]\sim\log N\frac{\partial^{2}}{\partial\xi^{2}}w(\xi)$ and the eigenmodes at the center are phononic waves [251]. This expression also reproduces the behavior of the axial dispersion relation of a periodic chain for $k\sim\pi/(Na)$, and $\omega^{\|}\sim k\sqrt{|\log k|}$ [8].[250]. In [182] it was experimentally shown that the micromotion modifies the eigenfrequencies close to the transition. The measurements are in agreement with a theoretical treatment, which includes the periodic drive of the Paul trap [198] (see also Sec. III.1).
@@
Figure 16: (Color online) (a) Low-frequency spectrum as a function of the aspect ratio $\alpha$. The solid lines are the modes of the harmonic crystal at T = 0, the symbols refer to the experimental measurements obtained with different techniques along with estimated uncertainties. The ions are laser cooled close to the Doppler limit. Comparison with molecular dynamics simulations gives $T\approx 3.5$ mK. (b) Schematic illustration of the Landau potential in the linear ($\alpha>\alpha_{c}$) and in the zigzag symmetry-broken phase ($\alpha<\alpha_{c}$). The coordinate $y$ indicates the transverse displacement of the central ion from the chain axis. Thermal excitations can overcome the barrier between the two degenerate zigzag configuration by inducing collective jumps of the crystal configuration. (c) Average dwelling time $\tau_{e}$, calculated using Kramers formula, in units of the zigzag mode period $T_{zz}=1/\omega_{k_{0}}$. In the simulations, temperatures $T$ are 0.1 mK (blue circles), 0.5 mK (orange squares), 2.0 mK (green diamonds), 3.5 mK (red triangles). The lines are a guide for the eye. For T=T=0.1 mK and 0.5 mK, missing points indicate no switches were observed during the simulation time. The dotted horizontal line indicates $\tau_{e}=T_{zz}$ . Inset shows the potential barrier $E_{B}$ for different trapping ratios $\alpha$. Adapted from Ref. 189.
@@
At finite, non-vanishing temperatures, the theory predicts that the linear-zigzag transition is a smooth crossover. For finite chains, nevertheless, a relatively sharp transition can be measured when the chain is prepared at temperatures below a critical value $T_{c}$ that depends on the number of ions $N$. In Ref. [189] the normal-mode spectrum of a trapped ion chain was measured at the symmetry-breaking linear to zigzag transition and at finite temperatures. Spectroscopy was performed by modulating the amplitude of the Doppler cooling laser to excite and measure mode oscillations of a chain composed of 30 172Yb ions. The expected mode softening at the critical point, a signature of the second-order transition, was not observed, see the left panel of Fig. 16. The spectroscopic signal at the transition is reproduced by molecular dynamics simulations, suggesting that the disappearance of the mode softening is due to the finite temperature of the chain. The experimental low-frequency spectrum about the transition point is also reproduced by an effective theory based on Landau model, where thermal effects determine the weight of anharmonicities coupling the soft mode with high-frequency modes [189]. Inspection of the numerical trajectories shows that the ions collectively jump between the two ground-state configurations of the symmetry-broken phase, as illustrated in the right panel of Fig. 16. This jump rate is qualitatively reproduced using Kramers’ formalism [81]. In this regime, where the ion dynamics consists of thermally activated jumps between the two zigzag configurations, a time-averaged measurement of the mean transverse displacement detects a linear chain where one would instead expect a zigzag structure. The transition to a zigzag is measured at trap aspect ratios below the critical one, resembling a temperature-induced shift of the transition point [135, 81, 210].
@@
In Ref. [70] a quantum thermodynamic study of the linear-zigzag transition was performed. Here, it is argued that the average work distribution diverges at the linear-zigzag critical point. It is proposed to reveal this prediction by performing small quenches of the ion trap frequency across the transition and measuring the ion transverse displacements distribution. Interferometric protocols, based on spin-motion entanglement across the linear-zigzag instability, could access the autocorrelation function of the crystal [74] and measure the normal mode spectrum at criticality [12, 13]. However, the interferometric signals are relatively sensitive to thermal effects, which tend to suppress entanglement [14]. So far, the universal critical exponents of Landau theory have been indirectly verified by experiments analyzing the scaling of defects by slowly quenching the trap aspect ratio across the linear-zigzag phase transition, see Sec. III.4. In Ref. [377] sideband spectroscopy was performed at the linear-zigzag transition, revealing quantum effects. This suggests a potential access to a structural transition driven by quantum fluctuations, that is discussed in the next subsection.
@@
Figure 17: (Color online) Left panel. Top: Sample images of the two symmetry broken equilibrium structures of the five 171Yb ion crystal far from the critical point. Bottom: Simulated quantum energy-level spectrum for the first two excited states of the transverse-$y$ zigzag mode of a four-ion crystal, calculated relative to the ground state. Color shading of the levels indicates the strength of Raman sideband coupling to the ground state for a central ion (see sidebar scales: right-handed bar for $0\rightarrow 1$ and left-handed bar for $0\rightarrow 2$). Also shown is the classical small-oscillation prediction and the form of the zigzag potential on either side of the transition. Right panel: measured frequency of first and second upper Raman sidebands for the transverse-y zigzag mode as a function of secular trap frequency $\nu_{y}$ relative to critical value $\nu_{y,c}=760$ kHz. Insets show the full range of data acquisition and a close-in view near the critical point. Solid lines are quantum energy-level differences of $n=1$ and $n=2$ number states with respect to the $n=0$ ground state for a quartic potential. Line shading corresponds to Raman coupling strength (see side-bar scale). Green dotted line shows the classical small-oscillation prediction for the full pseudopotential. From 377.
@@ II.1.5 The quantum linear-zigzag transition
Structural phase transitions in ion Coulomb crystals are typically modelled by the configurations that minimize the potential energy. Although this approach discards the kinetic energy, it proves to be sufficiently accurate to reproduce the experimental results for Doppler-cooled chains. It works even down to low dimensions, thanks to the long-range nature of the Coulomb interactions. For this reason, the predictions of Landau theory for the linear-zigzag transition are confirmed by several experimental results within the level of accuracy permitted by the temperature and by the measurement apparatus. The possibility to cool relatively long chains close to the zero-point motion, on the other hand, paves the way for verifying the effects of quantum fluctuations at the linear-zigzag transition. In the near vicinity of the transition, in fact, the competing influences of the Coulomb interaction and the confining potential balance each other out, and fluctuations are expected to become significant. In Ref. [293] tunnelling across the linear-zigzag transition was discussed for an anisotropic trap, which confines the motion of the ions on a plane. By means of a field-theoretical model, which was postulated on the basis of plausible assumptions, the tunneling rate between the two symmetry-broken zigzag ground states were estimated for chains of few ions. In Ref. [377] it was possible to experimentally reveal quantum tunnelling between zigzag configurations in a chain composed of three to five 171Yb ions and confined by a linear Paul trap. Figure 17 displays the spectroscopic signal of a Raman transition resolving the anharmonic spectrum of excitations of the zigzag mode, demonstrating the effect of quantum dynamics.
In Refs. [328, 329] a quantum field theoretical model was derived for an anisotropic potential as in Ref. [377], which pins the zigzag orientation in the xyx-y plane. The procedure permits to map the quantum linear-zigzag transition to an Ising model in transverse field. The derivation incorporates quantum fluctuations in the classical model, Eq. (45), and leads to a 1+1 dimensional quantum field theory described by the partition function $Z=\int\mathcal{D}\phi e^{-S[\phi]/\hbar}$ with the Euclidean action [328]
@@
$+\left.\frac{1}{2}K(\phi_{j}-\phi_{j+1})^{2}\right]\,$, (46)
@@
$H_{\mathrm{I}}=-\sum_{j=1}^{N}\left(J\hat{\sigma}_{j}^{z}\hat{\sigma}_{j+1}^{z}+h\hat{\sigma}_{j}^{x}\right)\,$, (47)
@@
$J=K\bar{\varrho}^{2}=K\frac{m}{4\mathcal{A}}\left(\nu_{t}^{(c)2}-\nu_{t}^{2}\right)\,$, (48)
@@
Figure 18: Phase diagram for a linear-zigzag transition, according to the mapping to the quantum Ising chain, where $T$ is the temperature and $\varepsilon$ is defined in (49). The quantum critical point, at $\varepsilon_{c}>0$, separates the linear from the zigzag phase at $T=0$. For $0<\varepsilon<\varepsilon_{c}$ quantum fluctuations dominate, and the crystal is in the linear (disordered) phase. The dashed lines indicate the boundaries of the quantum critical region. From 328.
Figure 19: Computed critical exponents and central charge of the action in Eq. (46). The theory values are the ones of the Ising model in transverse field, see e.g. 304. From 331.
The quantum theory predicts a critical point that is shifted from the mean field value $\nu_{t}=\nu_{t}^{(c)}$ ($\varepsilon=0$). The shift scales with the ratio between kinetic and Coulomb energy,
$\tilde{\hbar}=\frac{\hbar}{\sqrt{maQ^{2}\log 2}}\,$, (50)
@@
In Ref. [279] the shift $\varepsilon_{c}$ was computed as a function of the small dimensionless parameter, $\tilde{\hbar}$. This was perfomed by rewriting the partition function $Z=\int{\mathcal{D}}\phi e^{-S[\phi]/\hbar}$ in the form